# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v1

@@ 5.6. Specific Examples
CVE-2024-0244 — Out-of-bounds Write Mitigation Strategies value: “TP”, rationale: The response accurately reflects the mitigation strategies for CVE-2024-0244 as outlined in the context. It mentions updating the firmware, avoiding direct internet connections, using secure private networks, and employing various programming practices to prevent buffer overflows, all of which are supported by the context provided. provenance: response: To mitigate this vulnerability, it is recommended to update the firmware of the affected devices to versions beyond 03.07, as advised by Canon. Additionally, users should avoid direct internet connections and instead use a secure private network with a firewall. Other strategies include using programming languages that prevent buffer overflows, employing vetted libraries, implementing automatic buffer overflow detection mechanisms, and adhering to strict memory management practices. context: …we advise that our customers install the latest firmware available for the affected models… We also recommend that customers set a private IP address for their products and create a network environment with a firewall or Wired/Wi-Fi router that can restrict network access. Potential Mitigations… Use a language that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use a vetted library or framework that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, avd.aquasec.com, psirt.canon/advisory-information/cp2024-001, canon.jp/support/support-info/240205vulnerability-response, canon-europe.com/support/product-security-latest-news ——————– Metrics: Cosine Similarity: 0.7607, ROUGE-L: 0.2299
@@
The above example shows that ProveRAG-Aqua successfully provides an accurate (TP) response, with a rationale and substantiated provenance. Interestingly, both the response and a significant portion of the evidence were derived from the hyperlinks, specifically “canon.jp”. This underscores the critical importance of integrating a wide array of URLs to ensure a comprehensive coverage that enhances the system’s capability to provide well-supported evidence. The provided metrics correspond to the values shown in Figure 3. Next, we ask: “What if the LLM failed to provide the most accurate response?” Here is a ‘FN’ example:
@@
CVE-2024-0925 — Stack-based Buffer Overflow Mitigation Strategies value: “FN”, rationale: The response omits specific details about the potential mitigations that are present in the context. While it mentions general strategies like buffer overflow detection mechanisms and ASLR, it does not include the specific examples and additional context provided in the context, such as the specific compiler flags and techniques like StackGuard, ProPolice, and D3-SFCV. provenance: response: Potential mitigations include using automatic buffer overflow detection mechanisms, employing Address Space Layout Randomization (ASLR), and compiling software with features that randomize memory addresses to prevent predictable exploit paths. context: Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. Examples include: the Microsoft Visual Studio /GS flag, Fedora/Red Hat FORTIFY_SOURCE GCC flag, StackGuard, and ProPolice, which provide various mechanisms including canary-based detection and range/index checking. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, avd.aquasec.com, https://github.com/yaoyue123/iot/blob/main/Tenda/AC10U/ ——————– Metrics: Cosine Similarity: 0.7038, ROUGE-L: 0.1892
@@
6. Conclusion
@@
In this work, we focused on enhancing LLMs’ ability to analyze and respond to cyber threats by developing ProveRAG, a system that emulates an analyst’s approach to vulnerability analysis while self-critiques its own response. By integrating an automated summarizing retrieval technique of up-to-date web data, and a self-critique mechanism, ProveRAG alleviates the omission and hallucination problem of state-of-the-art LLMs. It generates comprehensive exploitation steps and mitigation strategies, backing each claim with evidence sourced from authoritative databases such as NVD, CWE, and Aqua. Our findings demonstrate that ProveRAG achieves exceptional accuracy, surpassing direct-prompting and chunking retrieval methods in vulnerability analysis. By providing verified evidence and detailed rationales, ProveRAG not only equips analysts with timely, accurate insights but also ensures that the processes are well-documented for subsequent audits.
Acknowledgements.
@@
Appendix A Summarizing Technique Retrieval
The following shows the prompt used in Retr. LLM to assess each URL’s relevancy, and, if relevant, the LLM will summarize the URL content specifically to address either the Exploitation or the Mitigation question.
Relevancy Prompts Exploitation Prompt You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary. Content: content Please follow the steps below: Step 1: Assess Relevancy - Does the content provide relevant information to describe how this [CVE-xxx-xxx] can be exploited? - Answer: [Yes/No] Step 2: Summarize Relevant Information - If the answer is ”Yes” in Step 1: - Summarize the content with step-by-step description to exploit this vulnerability. - If the answer is ”No” in Step 1: - Summary: NONE Mitigation Prompt You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary. Content: content Please follow the steps below: Step 1: Assess Relevancy - Does the content provide relevant information to describe the potential mitigation strategies for this vulnerability? - Answer: [Yes/No] Step 2: Summarize Relevant Information - If the answer is ”Yes” in Step 1: - Summarize the content to describe the potential mitigation strategies for this CVE. - If the answer is ”No” in Step 1: - Summary: NONE
Appendix B Evaluation Module Prompts
This following blocks give the preamble of the evaluation prompt used in Eval. LLM. It defines the Value, the Rationale, and the Provenance for the Eval. LLM.
Evaluation Module Value prompt - Description: This attribute represents the accuracy of the response based on the provided context about the CVE-ID. - Allowed Values: - ‘TP’ (True Positive): The response fully and accurately reflects the information in the context. - ‘FP’ (False Positive): The response contains inaccurate or incorrect information that are not supported by the context. - ‘FN’ (False Negative): The response omits information that is present in the context. - Requirements: - Must strictly be one of the allowed values: ‘TP’, ‘FP’, or ‘FN’. - The selection must be based on an objective comparison between the response and the context. - Must follow the guidelines for Provenance. Rationale prompt - Description: This attribute provides the reasoning behind the selected ”value” for the evaluation of the response.
Evaluation Module Provenance prompt - Detailed Instructions for Provenance based on the selected ”value”: - If ”value” is ‘TP’: 1. Carefully compare the response with the context. 2. Identify and extract the key segments from both the response and the context that align perfectly. 3. Format the provenance as follows: - ‘response: ”[Extracted segment from the response that matches the context]”.’ - ‘context: ”[Corresponding segment(s) from the context]”.’ - If ”value” is ‘FP’: 1. Review the response thoroughly to identify parts that are incorrect or not present in the context. 2. Format the provenance as follows: - ‘response: ”[Incorrect or unsupported segment from the response]”.’ - ‘context: ”[The closest matching part from the context, or indicate ‘No corresponding information in context’ if none exists]”.’ - If ”value” is ‘FN’: 1. Examine the context to find essential information that is present but omitted in the response. 2. Format the provenance as follows: - ‘response: ”[The entire response]”.’ - ‘context: ”[Relevant segment from the context that should have been matched]”.’ Guidelines: - In all cases, ensure the provenance provided is clear, concise, and directly supports the selected ”value”. - The provenance should be an exact match or a direct comparison between the response and the context, following the formatting rules strictly.
Appendix C Examples contrasting Prompt-only vs. ProveRAG
This section gives two additional examples; the first one is generated with Prompt-Only experiment, and the second is generated via ProveRAG-Aqua:
For the CVE-2024-0338 (the Prompt-Only experiment), the system generated a ”FP” prediction, clearly illustrating a case of hallucination by the LLM. This is expected, as the CVE was disclosed in 2024, after the training data cutoff of the gpt-4o-mini model (October 2023). The provided metrics correspond to the values shown in Figure 3, reflecting this prediction.
For the CVE-2024-26029 (the ProveRAG-Aqua experiment), it is a “TP” prediction for exploitation information. The rationale and provenance fully support the selected value of ProveRAG. High Rouge-L and Cosine Similarity scores indicate strong alignment between the generated response and the supporting evidence. These metrics correspond to the values presented in Figure 3.
CVE-2024-0338 — Classic Buffer Overflow Exploitation Information value: “FP”, rationale: The response inaccurately describes the exploitation of CVE-2024-0338. It mentions a general process of exploitation involving a flaw in the system, but it fails to specify that this particular vulnerability is a buffer overflow in XAMPP, which allows arbitrary code execution through a long file debug argument. The response does not accurately reflect the specific nature of the vulnerability as described in the context. provenance: response: CVE-2024-0338 is a vulnerability that allows attackers to exploit a flaw in the system, potentially leading to unauthorized access or data manipulation. context: A buffer overflow vulnerability has been found in XAMPP affecting version 8.2.4 and earlier. An attacker could execute arbitrary code through a long file debug argument that controls the Structured Exception Handler (SEH). ——————– Metrics: Cosine Similarity: 0.4075, ROUGE-L: 0.1695
CVE-2024-26029 — Improper Access Control Exploitation Information value: “TP”, rationale: The response accurately reflects the information provided in the context regarding CVE-2024-26029. It correctly identifies the vulnerability as an Improper Access Control issue affecting Adobe Experience Manager versions 6.5.20 and earlier, and describes the exploitation method and consequences in a manner consistent with the context. provenance: response: CVE-2024-26029 is an Improper Access Control vulnerability affecting Adobe Experience Manager versions 6.5.20 and earlier. This vulnerability allows attackers to bypass security measures and gain unauthorized access to sensitive information without requiring user interaction. context: Adobe Experience Manager versions 6.5.20 and earlier are affected by an Improper Access Control vulnerability that could result in a Security feature bypass. An attacker could leverage this vulnerability to bypass security measures and gain unauthorized access. Exploitation of this issue does not require user interaction. ——————– Retrieved URLs: nvd.nist.gov, avd.aquasec.com ——————– Metrics: Cosine Similarity: 0.8250, ROUGE-L: 0.5116

### v2

@@ 5.7 Specific Examples
CVE-2024-0244 — Out-of-bounds Write –> GPT Mitigation Strategies value: “TP”, rationale: The response accurately reflects the mitigation strategies for CVE-2024-0244 as outlined in the context. It mentions updating the firmware, avoiding direct internet connections, using secure private networks, and employing various programming practices to prevent buffer overflows, all of which are supported by the context provided. provenance: response: To mitigate this vulnerability, it is recommended to update the firmware of the affected devices to versions beyond 03.07, as advised by Canon. Additionally, users should avoid direct internet connections and instead use a secure private network with a firewall. Other strategies include using programming languages that prevent buffer overflows, employing vetted libraries, implementing automatic buffer overflow detection mechanisms, and adhering to strict memory management practices. context: …we advise that our customers install the latest firmware available for the affected models… We also recommend that customers set a private IP address for their products and create a network environment with a firewall or Wired/Wi-Fi router that can restrict network access. Potential Mitigations… Use a language that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use a vetted library or framework that does not allow this weakness to occur or provides constructs that make this weakness easier to avoid… Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, avd.aquasec.com, psirt.canon/advisory-information/cp2024-001, canon.jp/support/support-info/240205vulnerability-response, canon-europe.com/support/product-security-latest-news ——————– Metrics: Cosine Similarity: 0.7607, ROUGE-L: 0.2299
@@
The above example shows that ProveRAG-Aqua successfully provides an accurate (TP) response, with a rationale and substantiated provenance. Interestingly, both the response and a significant portion of the evidence were derived from the hyperlinks, specifically “canon.jp”. This underscores the critical importance of integrating a wide array of URLs to ensure a comprehensive coverage that enhances the system’s capability to provide well-supported evidence. The provided metrics correspond to the values shown in Figure 3. Here is an hallucination (FP) example of gpt-4o-mini:
CVE-2024-0338 — Classic Buffer Overflow –> GPT Exploitation Information value: “FP”, rationale: The response inaccurately describes the exploitation of CVE-2024-0338. It mentions a general process of exploitation involving a flaw in the system, but it fails to specify that this particular vulnerability is a buffer overflow in XAMPP, which allows arbitrary code execution through a long file debug argument. The response does not accurately reflect the specific nature of the vulnerability as described in the context. provenance: response: CVE-2024-0338 is a vulnerability that allows attackers to exploit a flaw in the system, potentially leading to unauthorized access or data manipulation. context: A buffer overflow vulnerability has been found in XAMPP affecting version 8.2.4 and earlier. An attacker could execute arbitrary code through a long file debug argument that controls the Structured Exception Handler (SEH). ——————– Metrics: Cosine Similarity: 0.4075, ROUGE-L: 0.1695
For the CVE-2024-0338 (the Prompt-only experiment), the system generated a “FP” prediction, clearly illustrating a case of hallucination by the LLM. This is expected, as the CVE was disclosed in 2024, after the training data cutoff of the gpt-4o-mini model (Oct. 2023). The provided metrics correspond to the values shown in Figure 3, reflecting this prediction. Next, we ask: “What if the LLM failed to provide the most accurate response?” Here is a ‘FN’ example with gpt-4o-mini:
@@
CVE-2024-0925 — Stack Buffer Overflow –> GPT Mitigation Strategies value: “FN”, rationale: The response omits specific details about the potential mitigations that are present in the context. While it mentions general strategies like buffer overflow detection mechanisms and ASLR, it does not include the specific examples and additional context provided in the context, such as the specific compiler flags and techniques like StackGuard, ProPolice, and D3-SFCV. provenance: response: Potential mitigations include using automatic buffer overflow detection mechanisms, employing Address Space Layout Randomization (ASLR), and compiling software with features that randomize memory addresses to prevent predictable exploit paths. context: Use automatic buffer overflow detection mechanisms that are offered by certain compilers or compiler extensions. Examples include: the Microsoft Visual Studio /GS flag, Fedora/Red Hat FORTIFY_SOURCE GCC flag, StackGuard, and ProPolice, which provide various mechanisms including canary-based detection and range/index checking. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, avd.aquasec.com, https://github.com/yaoyue123/iot/blob/main/Tenda/AC10U/ ——————– Metrics: Cosine Similarity: 0.7038, ROUGE-L: 0.1892
@@
Now, to see the performance of ProveRAG on other LLM (and provenance quality), here are the responses of CVE-2024-0267 for both Llama3.1-8B and gpt-4o-mini:
CVE-2024-0267 — SQL Injection –> Llama Exploitation Information value: “FN”, rationale: This CVE (2024-0267) is a known vulnerability, and it has been disclosed to the public. The associated identifier of this vulnerability is VDB-249823. provenance: response: A vulnerability classified as critical was found in Kashipara Hospital Management System up to 1.0. Affected by this vulnerability is an unknown functionality of the file login.php of the component Parameter Handler. context: New CVE Received from VulDB ——————– Retrieved URLs: nvd.nist.gov, vuldb.com/?ctiid.249823 ——————– Metrics: Cosine Similarity: 0.3070, ROUGE-L: 0.0000
CVE-2024-0267 — SQL Injection –> GPT Exploitation Information value: “TP”, rationale: The response accurately describes the critical vulnerability CVE-2024-0267 in the Kashipara Hospital Management System, specifically detailing the SQL injection through the ’email’ and ’password’ parameters in the login.php file. It correctly states that an attacker can exploit this vulnerability to gain unauthorized access to the system’s database, which aligns with the context provided. provenance: response: CVE-2024-0267 is a critical vulnerability in the Kashipara Hospital Management System that allows for SQL injection through the manipulation of the ’email’ and ’password’ parameters in the login.php file. An attacker can exploit this vulnerability by injecting malicious SQL code into these parameters, which can lead to unauthorized access to the system’s database. context: A vulnerability classified as critical was found in Kashipara Hospital Management System up to 1.0. This vulnerability is known as CVE-2024-0267. The manipulation of the argument email/password leads to sql injection. The attack can be launched remotely. ——————– Retrieved URLs: nvd.nist.gov, cwe.mitre.org, vuldb.com/?id.249823 ——————– Metrics: Cosine Similarity: 0.8785, ROUGE-L: 0.3265
For CVE-2024-0267 with Llama-3.1-8B, the prediction is classified as a “FN” for exploitation information. The rationale and provenance validate the selected value in ProveRAG. As can be seen in the context, the Llama model failed to properly extract the retrieved information. Low Rouge-L and Cosine Similarity scores indicate the difference between the generated response and the supporting evidence. This shows how LLMs with weak instruction-following abilities will perform for provenance, and hence, indicating the need for either a more powerful model or a careful engineered prompt tailored to this model’s specific characteristics. In contrast, the gpt-4o-mini model accurately identified the prediction as “TP” by providing evidence from the Retrieved URLs, in which have strong overlaps with the model’s response with high Cosine Similarity and Rouge-L scores. These metrics correspond to the values presented in Figure 3. This comparison demonstrates how different models may struggle to effectively utilize retrieved context, emphasizing the role of model selection and/or a tailored, engineered prompt for a specific model to maximize the benefits of ProveRAG in optimizing performance and reliability in vulnerability analysis.
6 Conclusion
@@
In this work, we focused on enhancing LLMs’ ability to analyze and respond to cyber threats by developing ProveRAG, a system that emulates an analyst’s approach to vulnerability analysis while self-critiques its own response. By integrating an automated summarizing retrieval technique of up-to-date web data, and a self-critique mechanism, ProveRAG alleviates the omission and hallucination problem of state-of-the-art LLMs. It generates comprehensive exploitation steps and mitigation strategies, backing each claim with evidence sourced from authoritative databases such as NVD, CWE, and Aqua. Our findings demonstrate that ProveRAG achieves exceptional accuracy, surpassing direct-prompting and chunking retrieval methods in vulnerability analysis. Furthermore, our results supported ProveRAG’s adaptability and effectiveness across different LLMs, with the emphasis on the importance of model selection and tailored prompt engineering to optimize performance. By providing verified evidence and detailed rationales, ProveRAG not only equips analysts with timely, accurate insights but also ensures that the processes are well-documented for subsequent audits.
7 Ethics considerations
We believe this research does not raise ethical concerns, as it neither uncovers new vulnerabilities nor facilitates exploitation. Instead, this work focuses on developing a provenance-based LLM framework to enhance the analysis of vulnerabilities that are not present in the LLM’s knowledge base. By incorporating provenance, the system ensures factual accuracy and transparency in its generations, which allows for self-critique and providing evidence-based insights. The proposed framework operates solely on open-domain data and is designed to support cybersecurity professionals by offering reliable, verifiable outputs for vulnerability analysis. It does not introduce or expose previously undisclosed vulnerabilities, maintaining ethical integrity throughout its application. Furthermore, this research contributes to the broader understanding of the role and limitations of LLMs in cybersecurity. Our findings provide actionable insights for professionals regarding the reasonable expectations and ethical use of LLMs in cybersecurity contexts.
Acknowledgments